Proposals for IRAM Telescopes

The deadline for submission of observing proposals on IRAM telescopes, both the interferometer and the 30m, is

13 September 2007, 17:00 CEST (UT + 2 hours)

The scheduling period extends from 1 December 2007 to 31 May 2008. Proposals should be submitted through our web–based submission facility. Instructions can be found on our web page at URL:

http://www.iram.fr/GENERAL/

submission/submission.html

This web page also provides a link to the Call for Proposals where detailed information on time estimates, special observing modes, technical information and references for both the IRAM interferometer and the IRAM 30m telescope can be found. The submission facility will be opened about three weeks before the proposal deadline. Proposal form pages and the 30m time estimator are available now.

Please avoid last minute submissions when the network could be congested. As an insurance against network congestion or failure, we still accept, in well justified cases, proposals submitted by:

– fax to number: (+33) 476 42 54 69 or by

– ordinary mail addressed to:

IRAM Scientific Secretariat, 300, rue de la piscine, F-38406 St. Martin d’Hères, France

Proposals sent by e–mail are not accepted. Color plots will be printed/copied in grey scale. If color is considered essential for the understanding of a specific figure, a respective remark should be added in the figure caption. The color version may then be consulted in the electronic proposal by the referees.

Soon after the deadline the IRAM Scientific Secretariat sends an acknowledgement of receipt to the Principal Investigator of each proposal correctly received, together with the proposal registration number. To avoid the allocation of several numbers for the same proposal, send in your proposal only once. Note that the web facility allows cancelation and modification of proposals before the deadline. The facility also allows to view the proposal in its final form as it appears after re–compilation at IRAM. We urge proposers to make use of this possibility as we always receive a number of corrupted proposals (figures missing, blank pages, etc.).

Valid proposals contain the official cover page, up to two pages of text describing the scientific aims, and up to two more pages of figures, tables, and references. Proposals should not exceed these 5 pages of scientific material. Except for the technical pages for the interferometer, longer proposals will be cut.

The cover page, in postscript or in L^{A}TEX format, and the L^{A}TEX style file proposal.sty may be obtained from the IRAM web pages^{1} at URL

../GENERAL/submission/proposal.html. In case

of problems, contact the secretary, Cathy Berjaud (e–mail: berjaud@iram.fr). Please, make sure that your proposals use the current form pages.

In all cases, indicate on the proposal cover page whether your proposal is (or is not) a resubmission of a previously rejected proposal or a continuation of a previously accepted interferometer or 30m proposal. We request that the proposers describe very briefly in the introductory paragraph (automatically generated header “Proposal history: ”) why the proposal is being resubmitted (e.g. improved scientific justification) or is proposed to be continued (e.g. last observations suffered from bad weather).

Do not use characters smaller than 11pt. This could render your proposal illegible when copied or faxed. If we notice any formal problems before the deadline, we will make an effort to contact the principal investigator and solve the problem together.

Applications for zero spacing observations have been simplified. If the need for complementary 30m observations is evident already at the time when the PdB interferometer proposal is prepared, just note this need on the interferometer proposal. A separate proposal for the 30m telescope is not required. The blank form for interferometer proposals contains a bullet, labelled “zero spacing” which should then be checked. The interferometer style file will prompt for an additional paragraph in which the scientific need for the zero spacings should be described. It is essential to give here all observational details, including size of map, sampling density and rms noise, spectral resolution, receiver configuration and time requested.

A mailing list has been set up for astronomers interested in being notified about the availability of a new Call for Proposals. A link to this mailing list is on the IRAM web page. The list presently includes all users of IRAM telescopes during the last 2 years with valid e–mail addresses. Please check that your email address in this list is correct, and point out the existence of this list to interested colleagues.

J.M. Winters & C. Thum

1 from here on we give only relative URL addresses. In the absolute address the leading two dots (..) have to be replaced by the address of one of our mirror sites: http://www.iram.fr or http://www.iram.es.

Travel funds for European astronomers

IRAM is one of the organizations participating in the RadioNet project, an initiative funded by the European Commission within the FP6 Programme to improve and encourage communication among astronomers of the European Community and associated countries. Transnational access (TNA) is the largest RadioNet programme and provides funding for travel expenses incurred by eligible users for carrying out their observations or reducing their data. As a partner of RadioNet, IRAM has some limited TNA funds to pay travel expenses for eligible users. Detailed information about user eligibility, TNA contacts, policies and travel claims for the IRAM 30m telescope and Plateau de Bure Interferometer can be found on the RadioNet home page at

http://www.radionet-eu.org.

Observers requesting TNA support will be asked to provide the necessary personal and professional information to IRAM. Funding through RadioNet should be acknowledged in publications resulting from TNA supported observations.

R. Neri & C. Thum

Call for Observing Proposals on the 30m Telescope

Summary

Proposals for three types of receivers will be considered for the coming winter semester:

1. the observatory’s set of four dual polarization heterodyne receivers centered at wavelengths of 3, 2, 1.3, and 1.1 mm.

2. the 9 pixel dual-polarization heterodyne receiver array, HERA, operating at 1.3 mm wavelength

3. The MAMBO–2 bolometer array with 117 pixels operating at 1.2 mm; the smaller MAMBO– 1 array with 37 pixels is kept as a backup.

Emphasis will be put on observations at the shorter wavelengths. About 2800 hours of observing time will be available.

The main news relevant for the coming winter semester are described below. The details of proposal formalities, instrumentation, observing modes, and estimation of observing time are described on the IRAM web site.

What is new?

Since November 2005, the telescope runs under a VME and Linux based new control system (NCS). Virtually all functionality of the old control system has been recovered. Notable exceptions are the coordinate system (only J2000 is currently supported) and the RASTER command. OTF, ONOFF or TRACKING (with frequency switching) commands together with SIC loops should be used instead. Observations with the rotated wobbler (non–azimuth wobbling directions) which are of interest mainly for bolometer mapping, are possible again. Other ongoing work concerns fine–tuning the tracking and slewing motions of the telescope.

Remote observing is available from the IRAM offices in Granada and Grenoble, and from the MPIfR in Bonn. The remote stations in Madrid and Paris are planned to become available also during this winter semester.

The NCS team maintains a web page (../IRAMES/ncs30m) where the current NCS status is described in detail.

The dual polarization HERA is operational together with its backends for high (VESPA) and low spectral resolution (WILMA, 4 MHz filters). Although tuning parameters are now available for a large range of frequencies, it is still requested to send us HERA frequencies in advance.

Like last semester, a bolometer array, preferentially the 117–channel MAMBO–2, will be available. Users should be aware that this instrument has lately shown some performance instabilities. Trouble shooting and maintenance efforts might reduce the available observing time.

A replacement of the current single pixel receivers by a modern more compact receiver is planned for next year. The new receiver will have strongly improved sensitivities and much larger IF bandwidth. However in a first time the frequency range below 83 GHz will not be covered. In view of this fact we encourage astronomers with an interest in frequencies below 83 GHz to propose such observations now.

The full Call for 30m Proposals is available on the IRAM web site.

Clemens Thum & Rainer Mauersberger